# Strongly Interacting Dark Matter admixed Neutron Stars

### v2

@@ 4 Results
In this section we show our results of the TOV-equation. We investigate the three mentioned equations of state for ordinary matter as well as two equations of state for dark matter described by one-flavor G2-QCD at two different values of the bare quark mass labelled as light and heavy. They correspond to the lightest particle, the dark pion, having a mass of 25% and 33% of our dark matter candidate, respectively Wellegehausen:2013cya . On top of that we investigate the mass $m_{C}$ of the dark matter candidate from 250 MeV to 4 GeV, and showing three benchmark values for $m_{C}$ in the figures. This range is well motivated from other studies on self-interacting dark matter and covers all interesting dynamics in the system Kulkarni:2022bvh ; Dengler:2024maq ; Hochberg:2014kqa . Also, fixing the scale is necessary as the combination of Newton’s constant and the mass of the neutron in ordinary matter sets an absolute scale. As will be seen, for dark matter masses much smaller or larger than the neutron the solutions can be obtained by scaling arguments. The mass and radius of a single-fluid compact object scale like $m^{-2}$ where $m$ is the scale with which the equation of state is rescaled. This scaling argument is recovered for very light dark matter, as it starts to dominate the neutron star. For heavy dark matter, we obtain an ordinary neutron star dominated by standard model matter. This is not too surprising as, despite all differences in details, the overall scales of G2-QCD are still similar to ordinary QCD Wellegehausen:2013cya . Thus, only if the ordinary matter scale and the dark matter scale are similar a balancing without either side dominating can be expected.
Figure 2: The investigated set of central pressures of ordinary matter ($p_{0,O}$) and dark matter ($p_{0,D}$). Gray points indicate stable solutions outside the constraints chosen in (12). Colored points indicate stable solutions that meet the constraints at >>1% dark matter fraction for the color coding and <<1% in yellow. The dark matter mass increases from 0.37 GeV (top) and 1 GeV (middle) to 4 GeV (bottom). The left-hand side shows the results for ordinary matter equation of state I, and the right hand side for III. Both are for the light dark matter equation of state.
For this intermediate mass range, we solved the two-fluid TOV equation by iterating over both the central pressures of ordinary matter and dark matter and perform the stability analysis. In figure 2 we show the range of central pressures that we investigated. Gray points indicate a stable solution to the TOV-equation. One immediate observation is that the addition of dark matter makes it possible to sustain larger central pressures for ordinary matter as was also found in Kain:2021hpk ; Valdez-Alvarado:2012rct . This is particularly interesting as this grants access to new regimes of equations of state that could potentially have interesting phenomenology. Moreover, heavier dark matter results in larger possible central pressures which can also be inferred from scaling arguments (see section 3.1).
As an additional constraint, we want to investigate under which circumstances the objects considered in this paper are compatible with astronomical observation. Given the recent developments, e. g. the observation of neutron star candidates with possibly surprising features Salmi:2024bss ; Doroshenko:2022nwp , we intentionally choose very relaxed bounds, to accommodate possible future discoveries. This will also allow our exploratory study to remain valid during possible future such observations. This leads to the following choice of bounds:
$\begin{split}0.7M_{\odot}<&M_{\text{tot}}<2.2M_{\odot}\\ 10\text{ km}<&R_{O}<16\text{ km}\\ &\Lambda<2000\end{split}$ (12)
where $M_{\text{tot}}=M_{O}+M_{D}$. Here, the bound on the radius only applies for the radius given by the ordinary matter as observation constraining the radius of neutron stars rely on electromagnetic radiation Ozel:2016oaf , which is expected to pass a potential dark matter halo unhindered. At the same time, mass is determined from gravitational effects, and will thus be provided by both components equally. In both cases the limits are chosen such as to loosely represent a conservative bound on astrophysically obtained result Ozel:2016oaf . Finally, the Love number is bounded from gravitational waves measurements LIGOScientific:2017vwq . We chose a more conservative bound to account for effects of dark matter and because the attribution of it to individual stars is difficult LIGOScientific:2017ync . The limit is therefore representing the most extreme possible cases. Again, this quantity is determined by a combination of dark matter and ordinary matter. The stable solutions to the TOV-equations that meet the criteria above are color coded by the amount of dark matter they contain $M_{D}/M_{\text{tot}}$. The maximal value of the color coding (45%) is given by the largest amount of dark matter consistent with the constraint within the three dark matter candidate masses shown. On top of that, we separate the cases where the neutron star contains only up to 1% dark matter mass and indicate these solutions by yellow points [^4] The color scheme of the points in this plot will be used throughout the paper.. Different publications use different realistic upper bounds for the amount of dark matter in neutron stars Ivanytskyi:2019wxd ; Miao:2022rqj ; Giangrandi:2022wht ; Rutherford:2022xeb . With our choice of 1% we want to ensure that the resulting compact objects can actually be identified with neutron stars. We will consider results that exceed this bound as non-standard compact stellar objects. As noted in the introduction, if the latter would be the dominating population, our reliance on a fixed QCD equation of state may become dubious.
Figure 3: The mass-radius relation between the observable ordinary radius $R_{O}$ in km and the total observable mass $M_{O}+M_{D}$ in solar masses in the same color-coding as figure 2. The left-hand side shows the results for ordinary matter equation of state I, and the right hand side for III. Both are for the light dark matter equation of state.
In figure 3 we show the mass radius relation for all stable neutron stars for 3 different values for the dark matter mass and for two of the three ordinary matter equations of state. These choices are representative. We thus relegate here and hereafter results for other setups to appendix 6.2. As expected, the impact of the dark matter is larger for light dark matter where we achieve very heavy objects that contain large amounts of dark matter. Although these are unlikely to be common neutron stars, they could potentially be interpreted as heavy non-standard compact stellar objects. This is an interesting possibility, as stiff equations of state are usually hard to obtain. By increasing the dark matter mass, its impact becomes more subtle in the sense that the resulting objects could very well be neutron stars with small dark matter parts.
Figure 4: The decomposition of the total mass in ordinary mass $M_{O}$ and dark mass $M_{D}$ in solar masses in the same color-coding as figure 2. The left-hand side shows the results for ordinary matter equation of state I, and the right hand side for III. Both are for the light dark matter equation of state. Note the different scales for $M_{D}$.
Figure 5: The observable radius of the ordinary matter $R_{O}$ in comparison to the radius of the dark matter $R_{D}$ in km in the same color-coding as figure 2. The left-hand side shows the results for ordinary matter equation of state I, and the right hand side for III. Both are for the light dark matter equation of state. The dotted line indicates where $R_{O}=R_{D}$. Solutions above indicate a dark matter halo scenario while solutions below indicate a dark matter core scenario.
We show contributions from the individual parts to the mass in figure 4 and for the radius in figure 5. We again also illustrate if the dark matter contribution to the total mass is limited to 1% of the total mass. The first important observation from figures 3-5 is that changing the dark matter mass by an order of magnitude has much more severe impact than changing the equation of state, except for the radius. The observable radius is basically entirely dominated by the ordinary matter equation of state.
The contribution of dark matter to the total mass is, on the other hand, primarily dominated by the mass of dark matter particles, with a decreasing contribution the larger the mass. The effect is stronger than linear, and can be traced back to the overall scaling of the equation of state of the dark matter by the dark matter mass. Still, once the dark matter mass is limited to at most 1% of the total mass, this effect becomes almost negligible. Note that very light dark matter allows for stable neutron stars with more than eight solar masses. While it is unlikely to be actually possible to stabilize the star at such masses, it indicates that the very rare observation of very heavy neutron stars could be linked to a dark matter component. Especially as very stiff equations of state for ordinary matter are not simple to obtain Friman:2011zz . In particular, our full dark matter equation of a QCD-like dark sector also does not in itself sustain very heavy neutron stars Hajizadeh:2017jsw . This may suggest that a separate investigation of the very heavy neutron stars, allowing for modifications of the ordinary matter equation of state, may be a worthwhile effort. This is an important hint, which could only be obtained from knowing a QCD-like equation of state, as is at our disposal.
The effective radius of the dark matter contribution is again strongly dominated by the dark matter mass, with the radius being inversely and non-linearly dependent on the dark matter mass. Especially for dark matter heavier or as heavy as a neutron, dark matter forms a core, while for very light dark matter a substantially extended dark matter halo is observable. In combination with the aforementioned results on the mass, this gives an interesting possibility. For light dark matter and heavy neutron stars, the dark matter halo in a neutron star merger should interact earlier than the ordinary matter core. That could give a modification of the gravitational wave signal Ivanytskyi:2019wxd . In combination with relaxing the stiffness of the ordinary matter equation of state, this would be a highly interesting future research project.
@@ 5 Summary
In this work we studied the impact of a strongly-interacting dark matter component on neutron star properties, using G2-QCD. The equation of state for the dark fluid was obtained at finite density from first principles using lattice field theory Maas:2012wr ; Wellegehausen:2013cya ; Hajizadeh:2017jsw . We used the two equations of state at our disposal, differing in their mass scales, and found that they have similar influence on the resulting properties of the mixed star. On the standard model side, we covered a range of equations of state that capture the current constraints on the mass, radius and tidal deformability. They were used in Kurkela:2014vha by interpolating between low and high densities using piecewise polytropes.
As is well known from investigations of the ordinary matter equations of state, they all result in different mass-radius-relations. The added dark matter has similar effects on all three of them. For each combination of equations of state we varied the mass of the dark matter candidate and the central pressure of the ordinary fluid and dark fluid and performed a stability analysis. The addition of dark matter, in general, allows for a larger possible central pressures which should be kept in mind when investigating equations of state Kain:2021hpk ; Valdez-Alvarado:2012rct . As expected, we find that light dark matter has a stronger effect on the properties of the neutron star and that by tuning the ratio of central pressures one is also able to find solutions that fulfill the experimental bounds. For some combination of the dark matter candidate mass and the ratio of central pressures we find results that do not resemble neutron stars ($M_{DM}$$M_{tot}$>1%) but could provide an explanation for non-standard compact stellar objects. An astrophysical observation of exotic compact objects also in gravitational wave experiments would motivate a more careful analysis of these objects. We leave this for future exploration.
It is noteworthy that, for the first time, this investigation used an equation of state from first principles for describing strongly-interacting dark matter in a two component neutron star. While the impact on, e. g., the mass-radius-relation may be similar to using an effective model, first principle calculations of a UV-complete confining theory that is also capable of addressing small-scale structure problems are an important step towards unveiling the true nature of dark matter.
Acknowledgements
@@
6 Appendix
6.1 Method of interpolation using polytropes
Figure 6: Labelling of the parameters of the piecewise polytropes
In order to use the lattice results for a calculation of the TOV equation we have to convert them to pressure $p$ and energy density $\varepsilon$. In order to ensure thermodynamic consistency at the level of guaranteeing $n\mu=p\varepsilon$, we use the framework of piecewise polytropes to interpolate between the data points
Following Read:2008iy , the polytropic equation of state is defined by Kurkela:2014vha
@@ 6.1 Method of interpolation using polytropes
$\mu=c+\frac{\Gamma}{\Gamma+1}Kn^{\Gamma-1}$ (15)
This formalism is used to interpolate between data points, see figure 6 for an illustration. The $i$’th piecewise polytrope, valid from $n_{i-1}$ to $n_{i}$, depends on the parameters $c_{i}$, $K_{i}$ and $\Gamma_{i}$. In the following, we will denote the functions depending on the number density and the $i$’th parameters as $p_{i}(n)$, $\varepsilon_{i}(n)$ and $\mu_{i}(n)$, and the values of these functions at $n_{i}$ as $p_{i}$, $\varepsilon_{i}$, $\mu_{i}$.
@@
In principle, this could again be cured, by combining more data points to fix the derivatives, introducing new parameters. This can be escalated, however this will always yield a discontinuity at latest at the $n$’th derivative for $n$ data points. This could only be avoided by a global fit, using a continuous function. This would, however, also eliminate phase transitions. Eventually, there is no optimal choice with having only a finite number of data points. Since the TOV equations do not depend on the speed of sound, and do not require it to be continuous, but only the (guaranteed) continuity and compatibility of the equation of state, polytropes are the minimal solution to guarantee this feature. The slight kinks are integrated over, and thus an over estimation and under estimation by the kinked solutions will average, to some degree, out. Hence, we choose this option here, following Kurkela:2014vha ; Dengler:2021qcq ; Pitz:2024xvh . The stability analysis in section 3.2 uses only integrated quantities and initial conditions, and is therefore not sensitive to local kinks in the equation of state.
We apply this framework for both the ordinary matter equations of state and the dark matter equation of state.
6.2 Results for the heavy dark matter equation of state
Figure 7: Same as figure 3, but for the heavy dark matter equation of state.
Figure 8: Same as figure 3, but the left column shows the results for the light dark matter equation of state and ordinary matter equation of state II, and the right column for the heavy dark matter equation of state and ordinary equation of state II.
In this appendix, we compile a set of results for the heavy equation of state for the dark matter sector, as well as for the equation of state II. We did not use them in the main text as they do not substantially change the overall picture. This can be seen for the mass-radius relation shown in figures 7 and 8, which need to be compared with figure 3.
6.3 Results for the Tidal Deformability
Figure 9: Results from the calculation of the tidal deformability for EoS I and the light ensemble in the same color-scheme as the plots in the main text. Left: The Love number $k_{2}$ plotted against the compactness. Right: The dimensionless tidal deformability $\Lambda$ plotted against the total mass $M_{tot}$.
For completeness, we show in figure 9 the results for the tidal deformability $\Lambda$, which is directly calculated from $k_{2}$ using eqn. 6 using the formalism described in section 3.3. We choose these equations of state to showcase the general behavior. We see similar features as in Dengler:2021qcq . The data points compatible with our constraints with a dark matter mass fraction <1% are essentially equal to the results for pure neutron star calculations. Allowing for a larger dark matter mass fraction, you obtain results with larger values of $k_{2}$ at smaller compactnesses. Given the scaling of $\Lambda$ with the compactness as $C^{-5}$, this will result in much larger values for $\Lambda$, which can also be seen on the right-hand side of the plot. We have already discussed the applicability of these results for gravitational wave signals in section 3.3. In the case shown here, the results close to a pure dark star yield slightly larger values for $k_{2}$ at larger values for C which in general results in smaller $\Lambda$. These kind of objects will become interesting if the merger of exotic compact objects is detected in a gravitational wave experiment.

### v3

@@ 4 Results
In this section, we show our results for the TOV equation. We investigated the three EoS of state for ordinary matter as well as two EoS for dark matter described by one-flavor G2-QCD at two different values of the bare quark mass labeled as light and heavy. They correspond to the lightest particle, the dark pion, having a mass of 25% and 33% of our dark matter candidate, respectively Wellegehausen et al. (2014). We will discuss the main findings by showing the results for EoS II and the light dark matter equation of state. The results for the remaining combinations of equations of state can be found in Section˜A.2, and differ only quantitatively. On top of that, we investigate the mass $m_{C}$ of the dark matter candidate from 500 MeV to 4 GeV, and show four benchmark values for $m_{C}$ in the figures. This range is well motivated by other studies on self-interacting dark matter and covers all interesting dynamics in the system Kulkarni et al. (2023); Dengler et al. (2024); Hochberg et al. (2015). Also, fixing the scale is necessary as the combination of Newton’s constant and the mass of the neutron in ordinary matter sets an absolute scale.
For dark matter masses much smaller or larger than the neutron the solutions can be obtained by scaling arguments. For a single fluid compact object whose EoS can be trivially scaled by some scale (the dark matter candidate mass in the case of our dark matter EoS), the mass and radius scale like $m^{-2}$ where $m$ is the scale. This scaling argument is recovered for very light dark matter, as it starts to dominate the neutron star. For heavy dark matter, we obtain an ordinary neutron star dominated by standard model matter. We expect a balance between the impact of the two fluids at similar mass scales, as, despite all differences in details, the overall scales of G2-QCD are still similar to ordinary QCD Wellegehausen et al. (2014).
Figure 2: The investigated set of central pressures of ordinary matter ($p_{0,O}$) and dark matter ($p_{0,D}$) for EoS II and the light dark matter equation of state. We will use this combination of ensembles in the main text and show further results in Section˜A.2. Gray points indicate stable solutions with a dark matter fraction >10%. Colored points indicate the dark matter fraction and we indicate solutions with dark matter fractions <<1% in yellow. The same color scheme is used in the following figures. We show four representative dark matter candidate masses from 0.5 GeV to 4 GeV.
We solve the two-fluid TOV equation by iterating over both the central pressures of ordinary matter and dark matter and perform the stability analysis. In Figure˜2 we show the range of central pressures that we investigated for EoS II and the light EoS for dark matter. The further results depend on the amount of dark matter in the object.
Different publications use different realistic upper bounds for the amount of dark matter in neutron stars Ivanytskyi et al. (2020); Miao et al. (2022); Giangrandi et al. (2023); Rutherford et al. (2023). Consequently, we will provide results not for a fixed amount of dark matter, but show the impact of a varying amount. In the figures, gray points indicate a stable solution to the TOV equation, but with a dark matter component larger than 10%. The further color coding indicates the dark matter fraction $M_{D}/M_{\text{tot}}$ below 10%. In particular, yellow points indicate solutions with dark matter fractions <1%. With our choice of 1% we want to ensure that the resulting compact objects can actually be identified with neutron stars. In general, for a dark matter component below 1% of the total mass, the results are more strongly affected by the choice of the ordinary matter EoS employed and than by the mass of the dark matter candidate.
One immediate observation is that the addition of (light) heavy dark matter makes it possible to sustain larger central pressures for (dark) ordinary matter as was also found in Kain (2021); Valdez-Alvarado et al. (2013). This is particularly interesting as this grants access to new regimes of EoS that could potentially have interesting phenomenology. We see that in the 4 GeV case, even when the dark matter fraction is kept at <1%, larger central pressures of ordinary matter are possible. We also find stable solutions that exceed this bound which could be labeled as non-standard compact stellar objects or dark stars. As noted in the introduction, if the latter would be the dominating population, our reliance on a fixed QCD EoS may become dubious.
Figure 3: The mass-radius relation between the total observable mass $M_{O}+M_{D}$ in solar masses and the observable ordinary radius $R_{O}$ in km in the same color-coding and for the same equations of state as in Figure˜2. The ellipses show results for the mass and radius of various neutron stars, which were determined using different techniques Doroshenko et al. (2022); Salmi and others (2024); Choudhury and others (2024); Abbott et al. (2017).
In Figure˜3 we show the mass-radius relation for all stable neutron stars using EoS II and the light equation of state. The y-axis is chosen to be the total mass, as mass measurements rely on gravitational signals. Radius measurements on the other hand, usually rely on electromagnetic observation, which is not sensitive to the dark matter substar. Therefore, the x-axis shows the ordinary matter radius. However, all configurations with a dark matter fraction >10% contain dark cores where the ordinary matter radius is identical to the outermost radius. We consistently observe that the mass decreases with the addition of dark matter. This is because the additional dark matter results in a larger central pressure which in turn increases the value of the metric function and with that a larger pressure gradient for both ordinary and dark matter. This results in a smaller core of the star, which is the densest part and therefore contributes the most to the mass. For all dark matter candidate masses, we find decent agreements with the observation of neutron stars, albeit that this statement mostly depends on the ordinary matter equation of state employed. However, we see that at 1% dark matter, the mass can shift by $\sim 0.1\,M_{\odot}$ and the radius by $\sim 100\,m$. For up to 10% dark matter, the mass can decrease by $\sim 0.4\,M_{\odot}$ and the radius can decrease by $\sim 800\,m$. The apparent observation that Doroshenko et al. (2022); Salmi and others (2024); Choudhury and others (2024); Abbott et al. (2017) are better described with a substantial dark matter content than with our pure ordinary matter EoS should not be interpreted too much. We do not model the ordinary matter crust, which especially could affect the apparent radius, and we do not fully sample all proposed ordinary matter EoS. What should be taken from these results, however, is that added SIMP dark matter can easily hide inside the observations. Especially, if, as is the case presently, it shows a similar, large, stiffness as required by the observations. This is an inherent feature of our dark matter EoS, and a direct consequence of the details of its self-interaction captured by the lattice simulations.
Figure 4: Left: The dark matter mass versus the total mass. Right: The dark matter radius versus the ordinary matter radius. The x-axes are chosen such that we can show the neutron star measurements from Doroshenko et al. (2022); Salmi and others (2024); Choudhury and others (2024); Abbott et al. (2017). Note the different y-axes for $R_{D}$.
Figure 5: The dimensionless tidal deformability versus the total mass. The yellow and red boxed are gravitational wave observation from Abbott et al. (2017) and Abbott and others (2020).
In Figure˜4, we show the dark matter mass versus the total mass and the dark matter radius versus the ordinary matter radius. The x-axes are chosen such that we can show neutron star measurements as horizontal bands. As we restrict ourselves to 10% dark matter in these plots, the total mass is mostly ordinary matter. In general, a light dark matter candidate will result in a larger dark matter extent. A heavy dark matter candidate on the other hand will result in a compact dark matter core. While the halo contributes more equally to the pressure gradient at all distances from the center of the star, the core has essentially no impact on the outer regions of the star. As a result, in the top panel on the left we see consistently an increase of the dark matter mass while the total mass decreases when the central pressure of dark matter is increased. In the bottom panel on the left, we see that increasing the dark matter central pressure first results in a larger total and dark matter mass. At some point, the dark core becomes large enough to influence the ordinary matter and results in a decreasing total mass at an increasing dark matter mass. On the right side, we see that the dark matter radius is much larger for small dark matter candidate masses. All configurations with a dark matter fraction <10% are core configurations meaning that $R_{D}<R_{O}$. For a 0.5 GeV dark matter candidate, the dark matter radius increases with increasing dark matter central pressure while the ordinary matter radius stays constant at first but decreases when the dark matter fraction exceeds 1%. For 4 GeV, we find that the dark matter radius increases only slightly by increasing the dark matter central pressure. The ordinary matter radius however decreases at first and increases afterwards roughly when the dark matter fraction exceeds 1%.
Finally, we discuss the results of dimensionless tidal deformability versus the total mass shown for the same configurations and the same color scheme in Figure˜5. Gravitational wave detection could be more promising for the detection of dark matter in neutron stars because it is also sensible to dynamical information. An extended halo will alter the gravitational wave signal in a different way from a dark matter core, and the latter could result in a supplementary peak in the gravitational wave spectrum. This has been shown in simulations of binary neutron star events Giangrandi et al. (2025). Our results cannot provide this dynamical information, but we can test our results with the constraints from neutron star merger observation. We find that when kept at <10%, the mixed stars are compatible with the measurements. In general, the addition of dark matter will decrease the tidal deformability but is barely visible on the logarithmic scale. We see that the extended halos for a 0.5 GeV dark matter candidate with dark matter fractions >10% will result in much larger tidal deformabilities, which is expected given the scaling with $R^{5}$. Mixed stars at larger dark matter candidate masses will decrease the tidal deformability.
@@ 5 Summary
In this work we studied the impact of a strongly-interacting dark matter component on neutron star properties, using G2-QCD. The EoS for the dark fluid was obtained at finite density from first principles using lattice field theory Maas et al. (2012); Wellegehausen et al. (2014); Hajizadeh and Maas (2017). We used the two EoS at our disposal, differing in their mass scales, and found that they have similar influence on the resulting properties of the mixed star. On the standard model side, we covered a range of equations of state that capture the current constraints on the mass, radius, and tidal deformability. They were used in Kurkela et al. (2014) by interpolating between low and high densities using piecewise polytropes.
As is well known from investigations of the ordinary matter EoS, they all result in different mass-radius-relations. The added dark matter has similar effects on all three of them. For each combination of EoS we varied the mass of the dark matter candidate and the central pressure of the ordinary fluid and dark fluid and performed a stability analysis. The addition of dark matter, in general, allows for larger possible central pressures which should be kept in mind when investigating ordinary matter EoS Kain (2021); Valdez-Alvarado et al. (2013). As expected, we find that light dark matter has a stronger effect on the properties of the neutron star and that by tuning the ratio of central pressures one is also able to find solutions that fulfill the observational bounds. For some combinations of the dark matter candidate mass and the ratio of central pressures we find mixed stars that do not resemble neutron stars but could provide an explanation for non-standard compact stellar objects, if encountered. An astrophysical observation of exotic compact objects also in gravitational wave experiments would motivate a more careful analysis of these objects. We leave this for future exploration.
It is noteworthy that, for the first time, this investigation used an EoS from first principles for describing strongly-interacting dark matter in a two component neutron star. In thus takes the impact of any self-interaction on the EoS fully and consistently into account, yielding e. g. its large stiffness uniquely. This is motivated by the class of SIMP dark matter models. While the impact on, e. g., the mass-radius-relation may be similar to using an effective model, first principle calculations of a UV-complete confining theory that is also capable of addressing small-scale structure problems are an important step towards unveiling the true nature of dark matter.
Acknowledgments
@@
Appendix A Appendix
A.1 Method of interpolation using polytropes
Figure 6: Labelling of the parameters of the piecewise polytropes.
In order to use the lattice results for a calculation of the TOV equation we have to convert them to pressure $p$ and energy density $\varepsilon$. In order to ensure thermodynamic consistency at the level of guaranteeing $n\mu=p+\varepsilon$, we use the framework of piecewise polytropes to interpolate between the data points
Following Read et al. (2009), the polytropic equation of state is defined by Kurkela et al. (2014)
@@ A.1 Method of interpolation using polytropes
$\mu=c+\frac{\Gamma}{\Gamma-1}Kn^{\Gamma-1}$ (14)
This formalism is used to interpolate between data points, see Figure˜6 for an illustration. The $i$’th piecewise polytrope, valid from $n_{i-1}$ to $n_{i}$, depends on the parameters $c_{i}$, $K_{i}$ and $\Gamma_{i}$. In the following, we will denote the functions depending on the number density and the $i$’th parameters as $p_{i}(n)$, $\varepsilon_{i}(n)$ and $\mu_{i}(n)$, and the values of these functions at $n_{i}$ as $p_{i}$, $\varepsilon_{i}$, $\mu_{i}$.
@@
In principle, this could again be cured, by combining more data points to fix the derivatives, introducing new parameters. This can be escalated, however this will always yield a discontinuity at latest at the $n$’th derivative for $n$ data points. This could only be avoided by a global fit, using a continuous function. This would, however, also eliminate phase transitions. Eventually, there is no optimal choice with having only a finite number of data points. Since the TOV equations do not depend on the speed of sound, and do not require it to be continuous, but only the (guaranteed) continuity and compatibility of the EoS, polytropes are the minimal solution to guarantee this feature. The slight kinks are integrated over, and thus an overestimation and underestimation by the kinked solutions will average, to some degree, out. Hence, we choose this option here, following Kurkela et al. (2014); Dengler et al. (2022); Pitz and Schaffner-Bielich (2024). The stability analysis in Section˜3.2 uses only integrated quantities and initial conditions, and is therefore not sensitive to local kinks in the EoS.
We apply this framework for both the ordinary matter EoS and the dark matter EoS.
A.2 Additional results
In this appendix, we show additional results for the masses, radii, and tidal deformabilities. Figure˜7 shows the mass-radius relations for the remaining combinations of EoS. First of all, we see that the shape of the mass-radius relation depends mostly on the ordinary matter EoS employed. This is expected as we restrict ourselves to <10% dark matter here. As a result, whether or not the mixed stars are compatible with the measurements also mostly depends on the ordinary matter EoS. We see, however, that the addition of dark matter can significantly change both the mass and the radius. We further see that the impact of the light dark matter and heavy dark matter EoS are very similar. We see a similar picture in the composition of the individual masses in in Figure˜8 but find that the light EoS can result in slightly higher dark matter masses. Figure˜9 shows the dark matter radius versus the ordinary matter radius. At low dark matter candidate masses, the impact of the light and heavy EoS are very similar. For large candidate masses, however, where the dark matter forms a compact core, the results for the light and heavy ensemble differ. The dark matter EoS results in much larger dark matter radii. Finally, in Figure˜10, we display the tidal deformability for the remaining combinations of EoS, which does not seem to be dependent on the dark matter EoS employed and mostly depends on the mass of the dark matter candidate.
In Figure˜11 we show the mass-radius relation with $m_{C}=0.5\,$GeV in a larger range to investigate solutions with large amounts of dark matter. The dark matter component in this case forms a large and massive dark matter halo which can reach a mass $4\,M_{\odot}$. These solutions do not resemble neutron stars at all, but could be interpreted as heavy non-standard compact stellar objects with very large masses at ordinary matter radii similar to neutron stars and smaller.
Figure 7: Same as Figure˜3, but for the remaining combinations of equations of state.
Figure 8: Same as the left of Figure˜4, but for the remaining combinations of equations of state.
Figure 9: Same as the right of Figure˜4, but for the remaining combinations of equations of state.
Figure 10: Same as Figure˜5, but for the remaining combinations of equations of state.
Figure 11: Same as Figure˜3, but for all combinations of equations of state using $m_{C}=0.5\,$GeV using a larger range to show the non-standard compact object solutions.